\begin{thebibliography}{99}

%\bibitem[AHY]{AHY} K.~Akagi, M.~Hirose, S.~Yasuda, {\itshape Integrality of $p$-adic multiple zeta values and application to finite multiple zeta values}, preprint.


%\bibitem[And99]{Andrews99} G.E.~Andrews, {\itshape $q$-analogs of the binomial coefficient congruences of Babage, Wolstenholme and Glaisher}, Discrete Math.~{\bf 204} (1999), 15--25.

%\bibitem{ApostolVu84} T.~M.~Apostol, T.~H.~Vu, {\itshape Dirichlet series related to the Riemann zeta function}, J. Number Theory {\bf 19} (1984) 85--102.

%\bibitem[AIK]{ArakawaIbukiyamaKaneko} T.~Arakawa, T.~Ibukiyama, M.~Kaneko, {\itshape Bernoulli Numbers and Zeta Functions}, Springer, Tokyo, 2014.

%\bibitem[BTT18]{BTT18} H.~Bachmann, Y.~Takeyama, K.~Tasaka, {\itshape Cyclotomic analogues of finite multiple zeta values}, Compositio Math. {\bf 154}(12) (2018), 2701--2721.

%\bibitem[BTT19]{BTT19} H.~Bachmann, Y.~Takeyama, K.~Tasaka, {\itshape Special values of finite multiple harmonic q-series at roots of unity}, IRMA Lectures in Mathematics and Theoretical Physics 31,
%Algebraic Combinatorics, Resurgence, Moulds and Applications (CARMA) (2020), Vol. 2,  1--18.
%IRMA Lectures in Mathematics and Theoretical Physics (EMS), CARMA Volume.

\bibitem{BachmannTakeyamaTasaka} H.~Bachmann, Y.~Takeyama, K.~Tasaka, {\itshape Finite and symmetric Mordell-Tornheim multiple zeta values}, to appear in J. Math. Soc. Japan.

%\bibitem[Bra05-1]{Bradley05} D.M.~Bradley, {\itshape Multiple $q$-zeta values}, J.~Algebra {\bf 283} (2005), no. 2, 752--798.

%\bibitem[Bra05-2]{Bradley05_2} D.M.~Bradley, {\itshape Duality for finite multiple harmonic $q$-series}, Discrete Math. {\bf 300} (2005), no. 1-3, 44--56.

\bibitem{BradleyZhou10} D.M.~Bradley, X.~Zhou, {\itshape On Mordell-Tornheim sums and multiple zeta values}, Ann.~Sci.~Math.~Qu\'{e}bec, {\bf 34} (2010), 15--23.

%\bibitem[Bro12]{Brown12} F.~Brown, {\itshape Mixed Tate motives over $\Z$}, Ann.~of Math.~{\bf 175} (2012), no.~2, 949--976.

%\bibitem{Chatzistamatiou17} A.~Chatzistamatiou, {\itshape On integrality of p-adic iterated integrals}, Journal of Algebra {\bf 474} (2017), 240--270.

%\bibitem{DeligneGoncharov05} P.~Deligne, A.~B.~Goncharov, {\itshape Groupes fondamentaux motiviques de Tate mixte}, Ann. Sci. \'{E}cole Norm. Sup. 38 (2005), 1--56.

%\bibitem{Dilcher08} K.~Dilcher, {\itshape Determinant expressions for q-harmonic congruences and degenerate Bernoulli numbers}, Electron.~J.~Combin.~{\bf 15}(1) (2008), R63.

%\bibitem{Drinfeld90} V.G.~Drinfeld, {\itshape On quasitriangular quasi-Hopf algebras and on a group that is closely connected with ${\rm Gal}(\overline{\Q}/\Q)$}, Algebra i Analiz 2 (1990), 149--181.

\bibitem{FurushoKomoriMatsumotoTsumura17} H.~Furusho, Y.~Komori, K.~Matsumoto, H.~Tsumura, {\itshape Desingularization of multiple zeta-functions of generalized Hurwitz-Lerch type and evaluation of $p$-adic multiple L-functions at arbitrary integers}. Various aspects of multiple zeta values, 27--66, RIMS Kokyuroku Bessatsu, {\bf B68}, Res. Inst. Math. Sci. (RIMS), Kyoto, 2017.

%\bibitem{Furusho04} H.~Furusho, {\itshape $p$-adic multiple zeta values. I. $p$-adic multiple polylogarithms and the $p$-adic KZ equation}, Invent.~Math., {\bf 155}, no. 2, (2004), 253--286.

%\bibitem{Furusho07} H.~Furusho, {\itshape $p$-adic multiple zeta values. II. Tannakian interpretations}, Amer.~J.~Math., {\bf 129}, no. 4, (2007), 1105--1144.



%\bibitem{FurushoJafari07} H.~Furusho, A.~Jafari, {\itshape Regularization and generalized double shuffle relations for $p$-adic multiple zeta values}, Compositio Math. Vol {\bf 143}, (2007), 1089--1107.

%\bibitem{Goncharov01} A.~Goncharov, \textit{Multiple polylogarithms and mixed Tate motives}, arXiv:math/0103059.

%\bibitem[GZ19]{GuoZudilin19} V.J.W.~Guo, W.~Zudilin, {\itshape A q-microscope for supercongruences}, Adv. in Math. {\bf 346} (2019) 329--358.

%\bibitem[Hir20]{Hirose} M.~Hirose, {\itshape Double shuffle relations for refined symmetric multiple zeta values}, Doc. Math. {\bf 25} (2020), 365--380

%\bibitem[HMO]{HiroseMuraharaOno} M.~Hirose, H.~Murahara, M.~Ono, {\itshape On variants of symmetric multiple zeta-star values and the cyclic sum formula}, to appear in Ramanujan J., arXiv:2001.03832.

%\bibitem[HHT17]{HPT17} Kh.~Hessami Pilehrood, T.~Hessami Pilehrood, R.~Tauraso, {\itshape Some $q$-congruences for homogeneous and quasi-homogeneous multiple $q$-harmonic sums}, Ramanujan J.~{\bf 43} (2017), 113--139.

%\bibitem[Hof97]{Hoffman97} M.E.~Hoffman, {\itshape The algebra of multiple harmonic series}, J.~Algebra {\bf 194} (1997), no. 2, 477--495.

%\bibitem[Hof15]{Hoffman15} M.E.~Hoffman, {\itshape Quasi-symmetric functions and mod $p$ multiple harmonic sums}, Kyushu J. Math. {\bf 69} (2015), 345--366.

%\bibitem{IharaKajikawaOhnoOkuda11} K.~Ihara, J.~Kajikawa, Y.~Ohno and J.~Okuda, {\itshape Multiple zeta values vs. multiple zeta-star values}, J.~Algebra {\bf 332} (2011), 187--208

%\bibitem[IKZ06]{IharaKanekoZagier06} K.~Ihara, M.~Kaneko and  D.~Zagier, {\itshape Derivation and double shuffle relations for multiple zeta values}, Compositio Math. {\bf 142} (2006), 307--338.


%\bibitem[Jar1]{Jarossay18} D.~Jarossay, {\itshape Pro-unipotent harmonic actions and a computation of $p$-adic cyclotomic multiple zeta values}, preprint (2018), arXiv:1501.04893v5


%\bibitem[Jar2]{Jarossay19} D.~Jarossay, {\itshape Adjoint cyclotomic multiple zeta values and cyclotomic multiple harmonic values}, preprint (2019), arXiv:1412.5099v5.

\bibitem{Kamano16} K.~Kamano, {\itshape Finite Mordell-Tornheim multiple zeta values}, Funct.~Approx.~Comment.~Math. {\bf 54} (2016), 65--72.




\bibitem{Kaneko19} M.~Kaneko, {\itshape  An introduction to classical and finite multiple zeta values}, Publications Math\'{e}matiques de Besan\c{c}on, no.1 (2019), 103--129.

\bibitem{KanekoZagier} M.~Kaneko and D.~Zagier, {\itshape Finite multiple zeta values}, in preparation.


%\bibitem[Kaw19]{Kawasaki} N.~Kawasaki, {\itshape Hyperlogarithms, Bernoulli polynomials, and related multiple zeta values}, Dissertation, Tohoku University (2019).

%\bibitem[KO20]{KawasakiOyama} N.~Kawasaki, K.~Oyama, {\itshape Cyclic sum of finite multiple zeta values}, Acta Arith. {\bf 195} (2020), 281--288.

%\bibitem[Kaw10]{Kawashima10} G.~Kawashima, {\itshape A generalization of the duality for finite multiple harmonic q-series}, Ramanujan J. {\bf 21} (2010), 335--347.

\bibitem{Komori08} Y.~Komori, {\itshape An integral representation of the Mordell-Tornheim double zeta function and its values at non-positive integers}, Ramanujan J., {\bf 17} (2008), 163--183.

\bibitem{Komori18} Y.~Komori, {\itshape Finite multiple zeta values, multiple zeta functions and multiple Bernoulli polynomials}, Kyushu J. Math. {\bf 72} (2018), 333--342.

\bibitem{Komori21} Y.~Komori, {\itshape Finite multiple zeta values, symmetric multiple zeta values and unified multiple zeta functions}, Tohoku Math.~J., {\bf 73}(2) (2021), 221--255.

%\bibitem{Kontsevich09} M.~Kontsevich, {\itshape Holonomic D-modules and positive characteristic}, Jpn. J. Math. {\bf 4} (2009), no. 1, 1--25.



		
\bibitem{Matsumoto02} K.~Matsumoto, {\itshape On the analytic continuation of of various multiple-zeta functions}, in “ Number Theory for the Millennium II, Proc. of the Millennial Conference on Number Theory”, M. A. Bennett et. al. (eds.), A. K. Peters, 2002, pp. 417--440.

\bibitem{Matsumoto03} K.~Matsumoto, {\itshape On Mordell-Tornheim and other multiple zeta-functions}, in Proceedings of the Session in Analytic Number Theory and Diophantine Equations, D. R. Heath-Brown and B. Z. Moroz (eds.), Bonner Math. Schriften 360, Bonn, 2003, n.25, 17pp.

\bibitem{MatsumotoNakamuraOchiaiTsumura08} K.~Matsumoto, T.~Nakamura, H.~Ochiai, H.~Tsumura, {\itshape On value-relations, functional relations and singularities of Mordell-Tornheim and related triple zeta-functions}, Acta Arith. {\bf 132} (2008), 99--125.

\bibitem{MatsumotoOnozukaWakabayashi19} K.~Matsumoto, T.~Onozuka, I.~Wakabayashi, {\itshape Laurent series expansions of multiple zeta-functions of Euler-Zagier type at integer points}, Math. Z. {\bf 295}, 623--642 (2020). 


\bibitem{Mordell58} L.~J.~Mordell, {\itshape On the evaluation of some multiple series}, J.~London Math.~Soc., {\bf 33} (1958), 368--371.


%\bibitem{Murahara16} H.~Murahara, \textit{A Note on Finite Real Multiple Zeta Values}, Kyushu J. Math. {\bf 70} (2016), no. 1, 197--204.

%\bibitem{Nakamura06} T.~Nakamura, {\itshape A functional relation for the Tornheim double zeta function}, Acta Arith. {\bf 125} (2006), 257--263.

\bibitem{Nakamura21} T.~Nakamura, {\itshape Symmetric Tornheim double zeta functions}, Abh. Math. Semin. Univ. Hambg. {\bf 91} (2021), 5--14. 

%\bibitem[OO07]{OhnoOkuda07} Y.~Ohno and J.~Okuda, {\itshape On the sum formula for the $q$-analogue of non-strict multiple zeta values}, Proc. Amer. Math. Soc. {\bf 135} (2007), no. 10, 3029--3037.

\bibitem{OkamotoOnozuka15} T.~Okamoto, T.~Onozuka, {\itshape Mean value theorems for the Mordell-Tornheim double zeta-function}, Ramanujan J., {\bf 37} (2015), 131--163.

%\bibitem{OkamotoOnozuka16} T.~Okamoto, T.~Onozuka, {\itshape Functional equation for the Mordell-Tornheim multiple zeta-function}, Funct. Approx. Comment. Math. {\bf 55}(2) (2016), 227--241.

%\bibitem{OkudaTakeyama07} J.~Okuda and Y.~Takeyama, \textit{On relations for the multiple $q$-zeta values}, Ramanujan J. {\bf14} (2007), no. 3, 379--387.

\bibitem{OnoSekiYamamoto21} M.~Ono, S.~Seki, S.~Yamamoto, {\itshape Truncated $t$-adic symmetric multiple zeta values and double shuffle relations}, Research in Number Theory {\bf 7} (2021), 15.

%\bibitem{Onodera11} K.~Onodera, {\itshape Generalized log sine integrals and the Mordell--Tornheim zeta values}, Trans.~Amer.~Math.~Soc. {\bf 363} (2011), 1463--1485.

%\bibitem{Pan08} H.~Pan, {\itshape A Generalization of Wolstenholme's Harmonic Series Congruence}, Rocky Mountain J.~Math., {\bf 38}(4) (2008), 1263--1269.

%\bibitem[PARI]{PARI} The PARI Group, {\itshape PARI/GP version 2.11.0}, Univ. Bordeaux, 2018, http://pari. math.u- bordeaux.fr/.

%\bibitem[Rac02]{Racinet02} G.~Racinet, {\itshape Doubles m\'{e}langes des polylogarithmes multiples aux racines de l\'{u}nit\'{e}}, Publ. Math. IHES 95 (2002), 185--231.

%\bibitem{Reu} C.~Reutenauer, {\itshape Free Lie algebras} (Oxford Science Publications, Oxford, 1993).

%\bibitem[Ros13]{RosenPHD} J.~Rosen, {\itshape The arithmetic of multiple harmonic sums}, Thesis (2013).

%\bibitem[Ros15]{Rosen15} J.~Rosen, {\itshape Asymptotic relations for truncated multiple zeta values}, J. Lond. Math. Soc. (2) {\bf 91} (2015), 554--572.

%\bibitem{Rosen19} J.~Rosen, {\itshape The completed finite period map and Galois theory of supercongruences}, Int. Math. Res. Not. {\bf 2019}, no. 23, 7379--7405.


%\bibitem{SakugawaSeki17} K.~Sakugawa and S.~Seki, {\itshape On functional equations of finite multiple polylogarithms}, J. Algebra {\bf 469} (2017), 323--357.

%\bibitem[Sek17]{SekiPHD} S.~Seki, {\itshape Finite multiple polylogarithms}, Doctoral dissertation in Osaka University, 2017.

%\bibitem[Sek19]{Seki19} S.~Seki, {\itshape The $\boldsymbol{p}$-adic duality for the finite star-multiple polylogarithms}, Tohoku Math. J. {\bf 71} (2019), 111--122.

%\bibitem{SaitoWakabayashi15} S.~Saito and N.~Wakabayashi, {\itshape Sum formula for finite multiple zeta values}, J.~Math.~Soc.~Japan {\bf 67} (2015), 1069--1076.

%\bibitem[SP07]{ShiPan07} L.-L.~Shi, H.~Pan, {\itshape A $q$-analogue of Wolstenholme's harmonic series congruence}, Amer. Math. Monthly, {\bf 114}(6) (2007), 529--531.

%\bibitem{Takeyama09} Y.~Takeyama, \textit{A $q$-analogue of non-strict multiple zeta values and basic hypergeometric series}, Proc. Amer. Math. Soc. {\bf 137} (2009), no. 9, 2997--3002.

%\bibitem{Takeyama12} Y.~Takeyama, \textit{Quadratic relations for a $q$-analogue of multiple zeta values}, Ramanujan J.  {\bf 27} (2012), no. 1, 15--28.

%\bibitem{Takeyama13} Y.~Takeyama, {\itshape The algebra of a $q$-analogue of multiple harmonic series}, SIGMA {\bf 9} (2013), 061, 1--15.

%\bibitem{Terasoma02} T.~Terasoma, \textit{Mixed Tate motives and multiple zeta values}, Invent. Math. 149 (2002), no. 2, 339--369.

\bibitem{Tornheim50} L.~Tornheim, {\itshape Harmonic double series}, Amer.~J.~Math., {\bf 72} (1950), 303--314.
		
		
%\bibitem{Tsumura02} H.~Tsumura, {\itshape On some combinatorial relations for Tornheim's double series}, Acta Arith. {\bf 105} (2002), 239--252.

%\bibitem{Tsumura05} H.~Tsumura, {\itshape On Mordell--Tornheim zeta values}, Proc.~Amer.~Math.~Soc., {\bf 133} (2005), 2387--2393.

%\bibitem{Tsumura07} H.~Tsumura, {\itshape On functional relations between the Mordell--Tornheim double zeta functions and the Riemann zeta function}, Math.~Proc.~Camb.~Phil.~Soc. {\bf 142} (2007), 395--405.


%\bibitem[Van96]{Hamme96} L.~Van Hamme, {\itshape Some conjectures concerning partial sums of generalized hypergeometric series}, $p$-adic functional analysis (Nijmegen, 1996) Lecture Notes in Pure and Appl. Math., vol. 192, Dekker, New York, 1997, pp. 223--236.

\bibitem{Witten91} E.~Witten, {\itshape On quantum gauge theories in two dimensions}, Commun.~Math.~Phys., {\bf 141} (1991), 153--209.

%\bibitem[Wol62]{Wolstenholme62} J.~Wolstenholme, {\itshape On certain properties of prime numbers}, Quart. J. Math. Oxford Ser. {\bf 5} (1862), 35--39.


%\bibitem{Yasuda16} S.~Yasuda, {\itshape Finite real multiple zeta values generate the whole space Z}, Int. J. Number Theory {\bf 12} (2016), no. 3, 787--812.

%\bibitem{Yasuda19} S.~Yasuda, {\itshape From $p$-adic multiple zeta values to finite multiple zeta values}, In: K.~Sakugawa, K.~Tasaka, Y.~Mishiba (eds.) Proceedings of 26th summer school on Number Theory "Multiple Zeta Value" (2018), 189--202.

\bibitem{Zagier94} D.~Zagier, {\itshape Values of zeta functions and their applications}, in First European Congress of Mathematics, Vol. II, A. Joseph et al. (eds.), Progr. Math. 120, Birkh\"{a}user, 1994, 497--512.

%\bibitem[Zha08]{Zhao08} J.~Zhao, {\itshape Wolstenholme type theorem for multiple harmonic sums}, Int. J. Number Theory, {\bf 4}(1) (2008), 73--106.

%\bibitem[Zha13]{Zhao13} J.~Zhao, {\itshape On $q$-analog of Wolstenholme type congruences for multiple harmonic sums}, Integers {\bf 13} (2013), A23.

%\bibitem[Zha16]{Zhao16} J.~Zhao, {\itshape Multiple zeta functions, multiple polylogarithms and their special values}, Series on Number Theory and Its Applications, World Scientific, Vol. {\bf 12} (2016).

%\bibitem[Zud19]{Zudilin19} W.~Zudilin, {\itshape Congruences for $q$-binomial coefficients}, Annals of Combinatorics {\bf 23} (2019), no. 3-4, 1123--1135.


\end{thebibliography}

\documentclass[dvipdfmx,12pt]{amsart}

\usepackage{amsmath,amsthm,amssymb,mathdots,shuffle}
\usepackage{setspace}
\usepackage{mathtools}
\usepackage{color}
\usepackage[capitalize]{cleveref}
\usepackage{autonum}
\allowdisplaybreaks
\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}

\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}

\numberwithin{equation}{section}

\newcommand{\Z}{{\mathbb Z}}

\newcommand{\C}{{\mathbb C}}
\newcommand{\Q}{{\mathbb Q}}

\newcommand{\ep}{{\varepsilon}}

\title{On a unified double zeta function of Mordell--Tornheim type}

\author[S.~Kadota]{Shin-ya Kadota}
\address[S.~Kadota]{Faculty of Fundamental Science\\
	National Institute of Technology (KOSEN), Niihama College\\
	Niihama-city, Ehime, 792-8580\\
	Japan}
\email{s.kadota@niihama-nct.ac.jp}

\author[T.~Okamoto]{Takuya Okamoto}
\address[T.~Okamoto]{
	Institute of Liberal Arts and Science\\
	Toyohashi University of Technology\\
	Tempaku-cho, Aichi, 441-8580\\
	Japan}
\email{okamoto.takuya.ze@tut.jp}

\author[M.~Ono]{Masataka Ono}
\address[M.~Ono]{
	Global Education Center\\
	Waseda University\\
	Shinjuku, Tokyo, 169-8050\\
	Japan}
\email{m-ono@aoni.waseda.jp}

\author[K.~Tasaka]{Koji Tasaka}
\address[K.~Tasaka]{Department of Information Science and Technology\\
	Aichi Prefectural University\\
	Nagakute-city, Aichi, 480-1198\\
	Japan
}
\email{tasaka@ist.aichi-pu.ac.jp}

\subjclass{11M32}

\keywords{Double zeta functions of Mordell--Tornheim type}

\begin{document}

\begin{abstract}
We consider {a} double zeta function of Mordell--Tornheim type and compute its values at non-positive integer points.
We then discuss a possible generalization of the Kaneko--Zagier conjecture for all integer points.
\end{abstract}

\maketitle


\section{Introduction}

The double zeta function of Mordell--Tornheim type is defined by 
\[ \zeta_{MT}(s_1,s_2;s_3)=\sum_{m,n\ge1} m^{-s_1}n^{-s_2}(m+n)^{-s_3},\]
which converges absolutely when $\Re (s_1+s_3),\Re (s_2+s_3)>1$ and $\Re (s_1+s_2+s_3)>2$ (see \cite[Theorem 2.2]{OkamotoOnozuka15}).
The special values of this function at positive integer points are first studied by Tornheim \cite{Tornheim50} and independently by Mordell \cite{Mordell58} for the case $s_1=s_2=s_3$, and also rediscovered by Witten \cite{Witten91} in his volume formula for certain moduli spaces related to theoretical physics (see also Zagier's number theoretical treatment \cite{Zagier94}).
As a function, Matsumoto \cite[Theorem 1]{Matsumoto02} proves that the function $\zeta_{MT}(s_1,s_2;s_3)$ can be analytically continued to the whole $\C^3$-space.
Its true singularities are also determined (see also \cite[Theorem 6.1]{MatsumotoNakamuraOchiaiTsumura08}).
As an application, it is clarified that non-positive integers $s_1,s_2,s_3$ are points of indeterminancy, i.e., the values of $\zeta_{MT}(s_1,s_2;s_3)$ at non-positive integer points depend on a limiting process.
Explicit formulas for these values in terms of generalized Bernoulli numbers are computed by Komori \cite[Theorem 3]{Komori08}.

In this paper, we wish to study special values of the {function}
\[\omega_{\mathcal{U}}(s_1,s_2,s_3)=(-1)^{ s_3}\zeta_{MT}(s_1,s_2;s_3)+(-1)^{ s_2}\zeta_{MT}(s_3,s_1;s_2)+(-1)^{ s_1}\zeta_{MT}(s_2,s_3;s_1),\]
where {$s_1,s_2,s_3$ are complex variables and set} $(-1)^s=e^{\pi i s}$.
The function $\omega_{\mathcal{U}}(s_1,s_2,s_3)$ originates from the previous work in \cite{BachmannTakeyamaTasaka}, where they introduce the values of $\omega_{\mathcal{U}}(s_1,s_2,s_3)$ at positive integer points as a Mordell--Tornheim type analogue of the symmetric multiple zeta values (see also \cite{OnoSekiYamamoto21}). 
One of main results of this paper is an explicit evaluation for {coordinatewise limits} of $\omega_{\mathcal{U}}(s_1,s_2,s_3)$ at non-positive integer points.

\begin{theorem}\label{thm:main}
For any non-negative integers $m_1,m_2,m_3\in \Z_{\ge0}$ {and $a,b,c\in\{1,2,3\}$ such that $\{a,b,c\}=\{1,2,3\}$}, we have
\[ {\lim_{\ep_a\rightarrow0}\lim_{\ep_b\rightarrow0}\lim_{\ep_c\rightarrow0}}\omega_{\mathcal{U}}(-m_1+\ep_{1},-m_2+\ep_{2},-m_3+\ep_{3}) = \begin{cases} 1 & (m_1,m_2,m_3)=(0,0,0)\\ 0 & (m_1,m_2,m_3)\neq (0,0,0)\end{cases} .\]
\end{theorem}

 The contents of this paper are as follows.
In order to motivate our work, we briefly recall the Kaneko--Zagier conjecture on finite/symmetric multiple zeta values in \S2.
\S3 gives a result on true singularities of the function $\omega_{\mathcal{U}}(s_1,s_2,s_3)$.
\S4 is devoted to proving Theorem \ref{thm:main}.
Finally, in \S5, we discuss a possible generalization of the Kaneko--Zagier conjecture for the multiple zeta function of Mordell--Tornheim type.

\section{Background}

This paper is to mimic the study of the unified multiple zeta function introduced by Komori \cite{Komori21}.
We briefly review his results to motivate our work.

For each $r\ge0$, we call a tuple $\boldsymbol{k}=(k_1,\ldots,k_r)$ of positive integers an index and $k_1+\cdots+k_r$ the weight.
We regard the empty index $\emptyset$ as the unique index of weight 0 (and $r=0$).
We set $F(\emptyset)$ to be a unit element for any function $F$ on indices.

We first recall the Kaneko--Zagier conjecture \cite{KanekoZagier}.
For each index $\boldsymbol{k}=(k_1,\ldots,k_r)$ and positive integer $n$, define the multiple harmonic sum $H_n(\boldsymbol{k})$ by
\[ H_n(\boldsymbol{k}) = \sum_{1\le n_1<\cdots<n_r\le n} \frac{1}{n_1^{k_1}\cdots n_r^{k_r}}.\]
The empty sum is understood as 0.
If the last component $k_r$ is greater than 1, the limit at $n\rightarrow \infty$ exists and is called the multiple zeta value, denoted by
\begin{equation}\label{eq:MZV}
\zeta( \boldsymbol{k})=\sum_{1\le n_1<\cdots<n_r} \frac{1}{n_1^{k_1}\cdots n_r^{k_r}}.
\end{equation}
Kaneko and Zagier introduce two different types of multiple zeta values constructed from the above objects.
One of them is the finite multiple zeta value $\zeta_{\mathcal{A}}(\boldsymbol{k})$ defined for each index $\boldsymbol{k}$ by
\[\zeta_{\mathcal{A}}(\boldsymbol{k})=\big(H_{p-1}(\boldsymbol{k})\mod p\big)_p\]
in the $\Q$-algebra $\mathcal{A}=\big(\prod_{p}\Z/p\Z\big) \big/ \big(\bigoplus_p \Z/p\Z\big)$, where $p$ runs over all primes.
Another is the symmetric multiple zeta value $\zeta_{\mathcal{S}}(\boldsymbol{k})$ defined for each index $\boldsymbol{k}$ by
\[ \zeta_{\mathcal{S}}(\boldsymbol{k}) = \sum_{a=0}^r (-1)^{k_{a+1}+\cdots+k_r}\zeta^\shuffle(k_1,\ldots,k_a)\zeta^\shuffle(k_r,\ldots,k_{a+1})\mod \pi^2\mathcal{Z}\]
in the $\Q$-algebra $\mathcal{Z}/\pi^2\mathcal{Z}$, where we denote by $\mathcal{Z}$ the $\Q$-algebra generated by all multiple zeta values and by $\zeta^\shuffle(k_1,\ldots,k_r)\in\mathcal{Z}$ the shuffle regularized multiple zeta value.
Let $\mathbb{I}^+$ be the set of all indices.
The main conjecture of Kaneko and Zagier states that for a finite subset $\{a_{\boldsymbol{k}}\in\Q\mid \boldsymbol{k}\in \mathbb{I}^+\}$ of $\Q$, we have
\begin{equation}\label{eq:KanekoZagierConjecture}
\sum_{\boldsymbol{k}\in \mathbb{I}^+} a_{\boldsymbol{k}} \zeta_{\mathcal{S}}(\boldsymbol{k})=0 \ \stackrel{?}{\Longleftrightarrow}  \ \sum_{\boldsymbol{k}\in \mathbb{I}^+} a_{\boldsymbol{k}} \zeta_{\mathcal{A}}(\boldsymbol{k})=0.
\end{equation}

In \cite{Komori21}, Komori defines the {function} $\zeta_{\mathcal{U}}(s_1,\ldots,s_r)$ by
\[ \zeta_{\mathcal{U}}(s_1,\ldots,s_r)=\sum_{a=0}^r (-1)^{s_{a+1}+\cdots+s_r}\zeta(s_1,\ldots,s_a)\zeta(s_r,\ldots,s_{a+1}).\]
It is a function analogue of symmetric multiple zeta values.
A crucial feature is that the function $\zeta_{\mathcal{U}}(s_1,\ldots,s_r)$ is entire.
Thus, one can study the values $\zeta_{\mathcal{U}}(\boldsymbol{k})\in \C$ for any tuple $\boldsymbol{k}$ of integers.
These values are conjecturally elements in the polynomial ring $\mathcal{Z}[\pi i]$ over $\mathcal{Z}$.
Indeed, for all index $\boldsymbol{k}\in \mathbb{I}^+$, Komori \cite[Theorem 1.4]{Komori21} shows that $\zeta_{\mathcal{U}}(\boldsymbol{k})\in \mathcal{Z}[\pi i]$, and also that 
\begin{equation}\label{eq:zeta_U_cong}
\zeta_{\mathcal{U}}(\boldsymbol{k})\equiv \zeta_{\mathcal{S}}(\boldsymbol{k}) \mod \pi i \mathcal{Z}[\pi i].
\end{equation}

Now, assuming the conjecture $\zeta_{\mathcal{U}}(\boldsymbol{k})\stackrel{?}{\in} \mathcal{Z}[\pi i]$, one extends the definition of the symmetric multiple zeta values $\zeta_{\mathcal{S}}(\boldsymbol{k})$ to any tuple $\boldsymbol{k}$ of integers by
\[\zeta_{\mathcal{S}}(\boldsymbol{k})=\zeta_{\mathcal{U}}(\boldsymbol{k})  \mod \pi i \mathcal{Z}[\pi i].\]
Hereafter, we denote by $\mathbb{I}$ the set of all tuple $(k_1,\ldots,k_r)\in \Z^r \ (r\ge0)$ of integers.
A natural question to ask is then whether the conjecture \eqref{eq:KanekoZagierConjecture} holds for all tuple of integers.
Namely, for a finite subset $\{a_{\boldsymbol{k}}\in\Q\mid \boldsymbol{k}\in \mathbb{I}\}$ of $\Q$, we have
\begin{equation}\label{eq:KomoriConjecture}
\sum_{\boldsymbol{k}\in \mathbb{I}} a_{\boldsymbol{k}} \zeta_{\mathcal{S}}(\boldsymbol{k})=0 \ \stackrel{?}{\Longleftrightarrow}  \ \sum_{\boldsymbol{k}\in \mathbb{I}} a_{\boldsymbol{k}} \zeta_{\mathcal{A}}(\boldsymbol{k})=0.
\end{equation}
Note that, since there is no convergence issue, the finite multiple zeta values $\zeta_{\mathcal{A}}(\boldsymbol{k})$ are already defined for all $\boldsymbol{k}\in \mathbb{I}$ and that they span the same space with the $\Q$-vector space generated by the set $\{\zeta_{\mathcal{A}}(\boldsymbol{k})\mid \boldsymbol{k}\in \mathbb{I}^+\}$ (see \cite{Kaneko19}).
The main result of Komori \cite[Theorem 1.5]{Komori21} states that the conjecture \eqref{eq:KomoriConjecture} holds when we restrict $\boldsymbol{k}$ to tuples of non-positive integers.
A partial evidence for depth 2 case is also given in \cite[Proposition 1.6]{Komori21}.

In this paper we wish to study the Kaneko--Zagier conjecture \eqref{eq:KanekoZagierConjecture} for multiple zeta values of Mordell--Tornheim type, which are independently established in \cite{BachmannTakeyamaTasaka} and \cite{OnoSekiYamamoto21}.
Our terminology follows \cite{BachmannTakeyamaTasaka}. 
In this analogy, for each $\boldsymbol{k}=(k_1,\ldots,k_r)\in \mathbb{I}^+$, let 
\[ \omega_n(\boldsymbol{k})=\sum_{\substack{n_1+\cdots+n_r=n\\n_1,\ldots,n_r\ge1}} \frac{1}{n_1^{k_1}\cdots n_r^{k_r}}\]
and define
\[ \omega_{\mathcal{A}}(\boldsymbol{k}) = \big( \omega_p(\boldsymbol{k}) \mod p\big)_p \in \mathcal{A}\]
for $r\neq1$.
We call this the finite multiple omega value for short.
Note that the finite multiple omega value is, up to sign, equal to the finite multiple zeta value of Mordell--Tornheim type, which is first studied by Kamano \cite{Kamano16}.
As a real counterpart of this, we define the symmetric multiple omega value $\omega_{\mathcal{S}}(\boldsymbol{k})$ for $\boldsymbol{k}=(k_1,\ldots,k_r)\in \mathbb{I}^+$ with $r\neq1$ by
\[\omega_{\mathcal{S}}(\boldsymbol{k})=\sum_{a=1}^r (-1)^{k_a}\zeta_{MT}(\underbrace{k_1,\ldots,k_{a-1}}_{a-1},\underbrace{k_{a+1},\ldots,k_r}_{r-a};k_a)\in \mathcal{Z}/\pi^2\mathcal{Z},\]
where 
\begin{equation}\label{eq:MT}
\zeta_{MT}(k_1,\ldots,k_{r-1};k_r)=\sum_{n_1,\ldots,n_{r-1}\ge1} \frac{1}{n_1^{k_1}\cdots n_{r-1}^{k_{r-1}}(n_1+\cdots+n_{r-1})^{k_r}}
\end{equation}
are the multiple zeta values of Mordell--Tornheim type which can be written as $\Q$-linear combinations of multiple zeta values \eqref{eq:MZV} (see \cite[Theorem 1.1]{BradleyZhou10}).
A similar object to $\omega_{\mathcal{S}}(\boldsymbol{k})$ is introduced by a different method in \cite[Definition 4.5]{OnoSekiYamamoto21} (which in the case $\mathcal{S}$ differs in a sign).
The main conjecture of \cite[Conjecture 1.5]{BachmannTakeyamaTasaka} (see also \cite[\S4]{OnoSekiYamamoto21}) is that for a finite subset $\{a_{\boldsymbol{k}}\in\Q\mid \boldsymbol{k}\in \mathbb{I}^+, r\neq1\}$ of $\Q$, we have
\begin{equation}\label{eq:BTTConjecture}
\sum_{\boldsymbol{k}\in \mathbb{I}^+} a_{\boldsymbol{k}} \omega_{\mathcal{S}}(\boldsymbol{k})=0 \ \stackrel{?}{\Longleftrightarrow}  \ \sum_{\boldsymbol{k}\in \mathbb{I}^+} a_{\boldsymbol{k}} \omega_{\mathcal{A}}(\boldsymbol{k})=0.
\end{equation}
It can be seen from \cite[Theorem 1.6]{BachmannTakeyamaTasaka} that the conjecture \eqref{eq:BTTConjecture} is supported by the Kaneko--Zagier conjecture \eqref{eq:KanekoZagierConjecture}.

An ultimate goal of this project is to establish a Mordell--Tornheim analogue of Komori's conjecture \eqref{eq:KomoriConjecture}.
For that purpose, we need to find a suitable {function} whose values at integer points modulo $\pi^2$ satisfy the same relations with the finite multiple omega values.  
As a natural candidate, we consider the following function:
\[\omega_{\mathcal{U}}(s_1,\ldots,s_r)=\sum_{a=1}^r (-1)^{s_a}\zeta_{MT}(\underbrace{s_1,\ldots,s_{a-1}}_{a-1},\underbrace{s_{a+1},\ldots,s_r}_{r-a};s_a).\]
Similarly to \eqref{eq:zeta_U_cong}, the congruence $\omega_{\mathcal{U}}(\boldsymbol{k})\equiv \omega_{\mathcal{S}}(\boldsymbol{k}) \mod \pi^2 \mathcal{Z}$ holds for all $\boldsymbol{k}\in \mathbb{I}^+$, since $\zeta_{MT}(\boldsymbol{k})$ converges absolutely for any $\boldsymbol{k}=(k_1,\ldots,k_r)\in \mathbb{I}^+$ with $r\neq1$.
As a function, we see from \cite[Theorem 1]{Matsumoto03} that the function $\omega_{\mathcal{U}}(s_1,\ldots,s_r)$ can be meromorphically continued to $\C^r$ and its possible singularities are on
\begin{align}
&s_{j_1}+s_{j_2}=1-l \quad (1\le j_1<j_2\le r ,\ l\in \Z_{\ge0}),\\
&s_{j_1}+s_{j_2}+s_{j_3}=2-l \quad (1\le j_1<j_2<j_3\le r ,\ l\in \Z_{\ge0}),\\
&\vdots\\
&s_{j_1}+\cdots+s_{j_{r-1}}=r-2-l \quad (1\le j_1<\cdots <j_{r-1}\le r ,\ l\in \Z_{\ge0}),\\
&s_1+\cdots+s_r=r-1.
\end{align}
While the {function} $\zeta_{\mathcal{U}}(s_1,\ldots,s_r)$ cancels all possible singularities which come from singularities of multiple zeta functions, some of the above singularities of $\omega_{\mathcal{U}}(s_1,\ldots,s_r)$ will be true.
Thus the generalization of the conjecture \eqref{eq:BTTConjecture} via the {function $\omega_{\mathcal{U}}(s_1,\ldots,s_r)$} is not straightforward.
In this paper, as a part of the future project, we explicate the situation for the case $r=3$ and then discuss a possible generalization of \eqref{eq:BTTConjecture} to all integer points.

\section{Singularities}

By \cite[Theorem 1]{Matsumoto02}, the true singularities of the function $\zeta_{MT}(s_1,s_2;s_3)$ lie on 
$s_1+s_3=1-l,\ s_2+s_3=1-l$ ($l\in\Z_{\ge0}$) and $s_1+s_2+s_3=2$. 
Hence the possible singularities of $\omega_{\mathcal{U}}(s_1,s_2,s_3)$ lie on the subsets of $\C^3$ defined by one of the equations:
\[ s_{a}+s_{b}=1-l\ \ (1\leq a<b\leq 3,\ l\in\Z_{\ge0}),\quad s_1+s_2+s_3=2.\]
We now prove that these are true singularities.

\begin{theorem}\label{thm:singularities}
All points on the subsets of $\C^3$ defined by one of the equations $s_{a}+s_{b}=1-l$\ \ $(1\leq a<b\leq 3,\ l\in\Z_{\ge0})$ and $s_1+s_2+s_3=2$ are true singularities of the function $\omega_{\mathcal{U}}(s_1,s_2,s_3)$.
\end{theorem}

\begin{proof}
We first recall the relevant material from \cite{Matsumoto02}.
For $s_1,s_2,s_3\in \C$, $M\in \Z_{>0}$ and $0<\eta<1$, let
\[I(s_1,s_2,s_3;M-\eta)=\frac{1}{2\pi i} \int_{(M-\eta)} \Gamma(s_3+z)\Gamma(-z)\zeta(s_1+s_3+z)\zeta(s_2-z)\,dz,\]
where the path of integration is the vertical line from $M-\eta-i\infty$ to $M-\eta+i\infty$.
From \cite[Eq.~(5.3)]{Matsumoto02}, we have
\begin{align}\label{MBI}
\begin{split}
\zeta_{MT}&(s_1,s_2;s_3)=\frac{\Gamma(s_2+s_3-1)\Gamma(1-s_2)}{\Gamma(s_3)}\zeta(s_1+s_2+s_3-1)\\
&\quad+\sum_{k=0}^{M-1}\binom{-s_3}{k}\zeta(s_1+s_3+k)\zeta(s_2-k)+\frac{1}{\Gamma(s_3)}I(s_1,s_2,s_3;M-\eta),
\end{split}
\end{align}
where $M$ is a sufficiently large positive integer. We find that the integral $I(s_1,s_2,s_3;M-\eta)$ is holomorphic on the region
$$
D_M=\{ (s_1,s_2,s_3)\in\C^3 \mid \ \Re s_3>-M+\eta, \Re (s_1+s_3)>1-M+\eta, \Re s_2<1+M-\eta\},
$$
because in $D_M$ the poles of the integrand are not on the path of integration. Since $M$ is arbitrarily large, by \eqref{MBI}, the function $\zeta_{MT}(s_1,s_2;s_3)$ can be meromorphically continued to $\C^3$. We note that, in the case $s_2=l$ for a positive integer $l$, the first and the second terms on the right-hand side of \eqref{MBI} are singular, where $M> l$, but these singularities cancel each other. Indeed, put $s_2=l+\ep$ for a positive integer $l$. 
Using the well-known asymptotic formula
\begin{equation}\label{eq:gamma}
\Gamma(s)=
\displaystyle\frac{(-1)^{-n}}{(-n)!}\frac{1}{s-n}+\gamma_{n}+O(|s-n|)\quad (n\in\Z_{\le0}),
\end{equation}
where $\gamma_n$ is the constant term of the expansion of $\Gamma(s)$ at $s=n \in \Z_{\le0}$, we have
\begin{align}
\zeta_{MT}(s_1,s_2;s_3)
&=\frac{\Gamma(s_3+l-1+\ep)\Gamma(1-l-\ep)}{\Gamma(s_3)}\zeta(s_1+s_3+l-1+\ep)\\
&\quad+\binom{-s_3}{l-1}\zeta(s_1+s_3+l-1)\zeta(1+\ep)+O(1)\\
&=\frac{\Gamma(s_3+l-1)}{\Gamma(s_3)}\frac{(-1)^{l-1}}{(l-1)!}\zeta(s_1+s_3+l-1)\frac{1}{\ep}\\
&\quad+\binom{-s_3}{l-1}\zeta(s_1+s_3+l-1)\frac{1}{\ep}+O(1)\\
&=O(1).
\end{align}

We now prove that $s_1+s_3=1-l \ (l\in \Z_{\ge0})$ determines the true singularity of the function $\omega_{\mathcal{U}}(s_1,s_2,s_3)$. 
Take $M\ge l+1$.
From \eqref{MBI} the singular part of $\zeta_{MT}(s_1,s_2;s_3)$ corresponding to $s_1+s_3=1-l$ comes from 
\[{\binom{-s_3}{l}}\zeta(s_1+s_3+l)\zeta(s_2-l).\]
{Since $\zeta_{MT}(s_2,s_3;s_1)=\zeta_{MT}(s_3,s_2;s_1)$}, the corresponding singular part of $\omega_{\mathcal{U}}(s_1,s_2,s_3)$ is
\begin{align}
(-1)^{ s_1}\binom{-s_1}{l}&\zeta(s_1+s_3+l)\zeta(s_2-l)+(-1)^{ s_3}\binom{-s_3}{l}\zeta(s_1+s_3+l)\zeta(s_2-l)\\
&=\left\{(-1)^{ s_1}\binom{-s_1}{l}+(-1)^{ s_3}\binom{-s_3}{l}\right\}\zeta(s_1+s_3+l)\zeta(s_2-l)\\
&=\left((-1)^{ s_1}-(-1)^{ -s_1}\right)\binom{-s_1}{l}\zeta(s_1+s_3+l)\zeta(s_2-l).
\end{align}
Since 
$$
\left((-1)^{ s_1}-(-1)^{ -s_1}\right)\binom{-s_1}{l}\zeta(s_2-l)\not\equiv 0,
$$
we see that $s_1+s_3=1-l$ determines the true singularity. 
Note that we have $\omega_{\mathcal{U}}(s_1,s_2,s_3)=\omega_{\mathcal{U}}(s_{\sigma(1)},s_{\sigma(2)},s_{\sigma(3)})$ for any permutation $\sigma\in \mathfrak{S}_3$.
Thus, $s_{a}+s_{b}=1-l\ \ (1\leq a<b\leq 3,\ l\in\Z_{\ge0})$ determine the true singularities.
 
For the case $s_1+s_2+s_3=2$, by \eqref{MBI} the corresponding singular part of $\omega_{\mathcal{U}}(s_1,s_2,s_3)$ comes from 
$$
\zeta(s_1+s_2+s_3-1)\times \sum_{j=0}^2 (-1)^{ s_{3+j}}\frac{\Gamma(s_{2+j}+s_{3+j}-1)\Gamma(1-s_{2+j})}{\Gamma(s_{3+j})},
$$
where we set $s_{a+j}=s_b$ if $a+j\equiv b \mod 3$ for $b=1,2,3$ and $a,j\in \Z$.
Hereafter, we abuse the same notation for other triplets of complex numbers or variables.
Since $s_2+s_3=2-s_1$, by Euler's reflection formula 
$$
\Gamma(s)\Gamma(1-s)=\frac{\pi}{\sin\pi s},
$$
we obtain 
\begin{align}
(-1)^{ s_{3}}\frac{\Gamma(s_{2}+s_{3}-1)\Gamma(1-s_{2})}{\Gamma(s_{3})} &=(-1)^{s_3}\frac{\Gamma(1-s_1)\Gamma(1-s_{2})}{\Gamma(s_{3})} \\
& = \frac{\pi^2}{\Gamma(s_1)\Gamma(s_2)\Gamma(s_3)} \frac{(-1)^{s_3}}{\sin \pi s_1 \sin \pi s_2}\\
&= \frac{\pi^2}{\prod_{a=1}^3 \Gamma(s_a)\sin \pi s_a } (-1)^{s_3}\sin \pi s_3\\
&= \frac{\pi^2}{\prod_{a=1}^3 \Gamma(s_a)\sin \pi s_a }\frac{(-1)^{2s_3}-1}{2i}.
\end{align}
Since 
$$
\frac{\pi^2}{2i\prod_{a=1}^3 \Gamma(s_a)\sin \pi s_a }\sum_{j=0}^2 \big( (-1)^{2s_{3+j}}-1\big)  \not\equiv 0
$$
holds for the case $s_1+s_2+s_3=2$, we conclude that $s_1+s_2+s_3=2$ determines the true singularity of $\omega_{\mathcal{U}}(s_1,s_2,s_3)$, which completes the proof.
\end{proof}

Our proof of Theorem \ref{thm:singularities} shows that if $(k_1,k_2,k_3)\in \Z^3$ is a singular point of the function $\omega_{\mathcal{U}}(s_1,s_2,s_3)$, then it is a point of indeterminancy.

\begin{remark}
Desingularizations of the double zeta function of Mordell--Tornheim type are studied in \cite{FurushoKomoriMatsumotoTsumura17,Nakamura21}.
\end{remark}

\section{Proof of Theorem \ref{thm:main}}

In this section, we first study an asymptotic behavior of the function $\zeta_{MT}(s_1,s_2;s_3)$ at $(s_1,s_2,s_3)=(-m_1,-m_2,-m_3)$ for non-negative integers $m_1, m_2, m_3$, and then, give a proof of Theorem \ref{thm:main}.

For $m_1,m_2,m_3\in \Z_{\ge0}$, let $M=2+m_1+m_3$.
Then $(-m_1,-m_2,-m_3)\in D_{M}$, so the integral $I(s_1,s_2,s_3;M-\eta)$ is analytic at $(-m_1,-m_2,-m_3)$. Hence we have the estimation
\begin{equation}\label{eq:I}
 \frac{1}{\Gamma(s_3)}I(s_1,s_2,s_3;M-\eta)= O(|s_3+m_3|).
\end{equation}
Note that a similar argument can be found in \cite[\S5]{MatsumotoOnozukaWakabayashi19}.

We now give the asymptotics of the function $\zeta_{MT}(s_1,s_2;s_3)$.
Let $B_k$ be the $k$-th Seki--Bernoulli number defined by
\[
\frac{te^t}{e^t-1}=\sum_{k\ge0}B_k\frac{t^k}{k!}.
\]
For non-negative integers $m_1, m_2, m_3$, let $m=m_1+m_2+m_3+2$ and define rational numbers $b(m_1,m_2;m_3)$ and $c(m_1,m_2;m_3)$ by
\begin{align}
b(m_1, m_2; m_3)&=m_1!m_2!m_3!\binom{m-1}{m_3}\frac{B_{m}}{m!},\\
c(m_1, m_2; m_3)&=\sum_{\substack{n_1+n_2=m_3\\n_1, n_2\ge0}}\binom{m_3}{n_1}\frac{B_{m_1+n_1+1}}{m_1+n_1+1}\frac{B_{m_2+n_2+1}}{m_2+n_2+1}.
\end{align}

\begin{lemma}\label{lem:MT_asymp}
For $m_1,m_2,m_3\in \Z_{\ge0}$ {and $a,b,c\in\{1,2,3\}$ such that $\{a,b,c\}=\{1,2,3\}$, we have
\begin{align}
&\lim_{\varepsilon_a\rightarrow 0}\lim_{\varepsilon_b\rightarrow 0}\lim_{\varepsilon_c\rightarrow 0}\zeta_{MT}(-m_1+\varepsilon_1,-m_2+\varepsilon_2;-m_3+\varepsilon_3)\\
&=\lim_{\varepsilon_a\rightarrow 0}\lim_{\varepsilon_b\rightarrow 0}\lim_{\varepsilon_c\rightarrow 0}\left\{ (-1)^{m_2}b(m_2, m_3; m_1)\frac{\varepsilon_3}{\varepsilon_2+\varepsilon_3}+(-1)^{m_1}b(m_3, m_1; m_2)\frac{\varepsilon_3}{\varepsilon_1+\varepsilon_3}+c(m_1, m_2; m_3)\right\}.
\end{align}}
\end{lemma}

\begin{proof}
Let $m=m_1+m_2+m_3+2$.
For the first term on the right side of \eqref{MBI}, we use \eqref{eq:gamma} and the expansion $\Gamma(s)=(n-1)!+O(|s-n|)$ at $s=n\in\Z_{\ge1}$ to obtain 
\begin{align}
&\frac{\Gamma(-m_2-m_3-1+\ep_2+\ep_3)\Gamma(1+m_2-\ep_2)}{\Gamma(-m_3+\ep_3)}\zeta(1-m+\ep_1+\ep_2+\ep_3)\\
&=\left\{(-1)^{m_3}m_3!\ep_3+O(|\ep_3|^2)\right\}\left\{\frac{(-1)^{m_2+m_3+1}}{(m_2+m_3+1)!}\frac{1}{\ep_2+\ep_3}+\gamma_{-m_2-m_3-1}+O(|\ep_2+\ep_3|)\right\}\\
&\quad\times\left\{m_2!+O(|\ep_2|)\right\}\left\{\zeta(1-m)+O(|\ep_1+\ep_2+\ep_3|)\right\}\\
&=(-1)^{m_2}b(m_2, m_3; m_1)\frac{\ep_3}{\ep_2+\ep_3}\\
&\quad+O\left(\left|\frac{\ep_2\ep_3}{\ep_2+\ep_3}\right|\right)+O\left(\left|\frac{\ep_3(\ep_1+\ep_2+\ep_3)}{\ep_2+\ep_3}\right|\right)+O(|\ep_3|)+O\left(\left|\frac{\ep_3^2}{\ep_2+\ep_3}\right|\right)\\
\end{align}
Here, for the last equality we have used $\zeta(1-m)=-B_m/m \ (m\in \Z_{\ge1})$.

For the second term on the right side of \eqref{MBI}, let $M=2+m_1+m_3$.
Since 
\begin{equation}\label{eq:binomial}
\binom{m-\ep}{k}=
\begin{cases}
\displaystyle\binom{m}{k}+O(|\ep|)&(m\ge k),\\
\displaystyle\frac{(-1)^{k-m}m!(k-m-1)!}{k!}\ep+O(|\ep|^2)&(m<k)
\end{cases}
\end{equation}
holds for $k, m\in\Z_{\ge0}$, we have
\begin{align}
&\sum_{k=0}^{M-1}\binom{m_3-\ep_3}{k}\zeta(-m_1-m_3+k+\ep_1+\ep_3)\zeta(-m_2-k+\ep_2)\\
&=\sum_{k=0}^{m_1+m_3}\binom{m_3-\ep_3}{k}\zeta(-m_1-m_3+k+\ep_1+\ep_3)\zeta(-m_2-k+\ep_2)\\
&\quad+\binom{m_3-\ep_3}{m_1+m_3+1}\zeta(1+\ep_1+\ep_3)\zeta(1-m+\ep_2)\\
&=c(m_1, m_2; m_3)+(-1)^{m_1}b(m_3, m_1; m_2)\frac{\ep_3}{\ep_1+\ep_3}\\
&\quad+O(|\ep_3|)+O(|\ep_1+\ep_3|)+O(|\ep_2|)+O\left(\left|\frac{\ep_3^2}{\ep_1+\ep_3}\right|\right)+O\left(\left|\frac{\ep_2\ep_3}{\ep_1+\ep_3}\right|\right).
\end{align}
Here, for the last equality we have also used the Laurent expansion of $\zeta(s)$ at $s=1$. Thus by \eqref{eq:I}, we obtain the desired result.
\end{proof}

In order to prove Theorem \ref{thm:main}, we use generating functions.
Let
\begin{align}
B(t_1, t_2, t_3)&=\sum_{m_1, m_2, m_3\ge0}(-1)^{m_1+m_2}b(m_1, m_2; m_3)\frac{t_1^{m_1}t_2^{m_2}t_3^{m_3}}{m_1!m_2!m_3!},\\
C(t_1, t_2, t_3)&=\sum_{m_1, m_2, m_3\ge0}(-1)^{m_3}c(m_1, m_2; m_3)\frac{t_1^{m_1}t_2^{m_2}t_3^{m_3}}{m_1!m_2!m_3!}.
\end{align}

\begin{lemma}\label{lem:gen}
We have
\begin{equation}\label{eq:gen_func} 
\sum_{j=0}^2 \big(B(t_{1+j}, t_{2+j}, t_{3+j})+C(t_{1+j}, t_{2+j}, t_{3+j})\big)=1,
\end{equation}
where we again put $t_{a+j}=t_b$ if $a+j\equiv b \mod 3$ for $b=1,2,3$ and $a,j\in \Z$.
\end{lemma}
\begin{proof}
Set
\[\beta(t)=\sum_{k\ge0}\frac{B_{k+1}}{(k+1)!}t^k=\frac{e^t}{e^t-1}-\frac{1}{t}.\]
One computes
\begin{align}
B(t_1, t_2, t_3)&=\sum_{ l, m_3\ge0}\left(\sum_{\substack{m_1+m_2=l\\m_1, m_2\ge0}}t_1^{m_1}t_2^{m_2}\right)(-1)^l\binom{l+m_3+1}{m_3}\frac{B_{l+m_3+2}}{(l+m_3+2)!}t_3^{m_3}\\
&=\sum_{l, m_3\ge0}\frac{t_1^{l+1}-t_2^{l+1}}{t_1-t_2}(-1)^l\binom{l+m_3+1}{m_3}\frac{B_{l+m_3+2}}{(l+m_3+2)!}t_3^{m_3}\\
&=\frac{1}{t_1-t_2}\sum_{k\ge1}\left(\sum_{\substack{l+m_3+1=k\\l, m_3\ge0}}(-1)^{l+1}\binom{k}{m_3}(t_2^{l+1}-t_1^{l+1})t_3^{m_3}\right)\frac{B_{k+1}}{(k+1)!}\\
&=\frac{1}{t_1-t_2}\sum_{k\ge1}\left((t_3-t_2)^k-t_3^k - (t_3-t_1)^k+t_3^k\right)\frac{B_{k+1}}{(k+1)!}\\
&=\frac{\beta(t_3-t_2)-\beta(t_3-t_1)}{t_1-t_2},
\end{align}
and
\begin{align}
C(t_1, t_2, t_3)&=\sum_{m_1, m_2, n_1, n_2\ge0}\frac{B_{m_1+n_1+1}}{m_1+n_1+1}\frac{B_{m_2+n_2+1}}{m_2+n_2+1}\frac{t_1^{m_1}t_2^{m_2}(-t_3)^{n_1+n_2}}{m_1!m_2!n_1!n_2!}\\
&=\prod_{j=1}^2 \left(\sum_{m_j, n_j\ge0}\frac{B_{m_j+n_j+1}}{(m_j+n_j+1)!} \binom{m_j+n_j}{m_j}t_j^{m_j}(-t_3)^{n_j}\right)\\
&=\beta(t_1-t_3)\beta(t_2-t_3).
\end{align}
Substituting these expressions into the left side of \eqref{eq:gen_func} and then using
\[ \sum_{j=0}^2 \frac{1}{(t_{1+j}-t_{2+j})(t_{3+j}-t_{2+j})} =0\quad \mbox{and} \quad \sum_{j=0}^2 \frac{e^{t_{1+j}-t_{3+j}}}{e^{t_{1+j}-t_{3+j}}-1} \frac{e^{t_{2+j}-t_{3+j}}}{e^{t_{2+j}-t_{3+j}}-1} =1, \]
we get the desired result.
\end{proof}

We are now in a position to prove Theorem \ref{thm:main}.

\begin{proof}[Proof of Theorem \ref{thm:main}]
By Lemma \ref{lem:MT_asymp}, we have
\begin{align}
&{\lim_{\ep_a\rightarrow0}\lim_{\ep_b\rightarrow0}\lim_{\ep_c\rightarrow0}}\omega_{\mathcal{U}}(-m_1+\ep_1, -m_2+\ep_2, -m_3+\ep_3)\\
&=\sum_{j=0}^2 (-1)^{m_{1+j}+m_{2+j}}b(m_{1+j}, m_{2+j}; m_{3+j}){\lim_{\ep_a\rightarrow0}\lim_{\ep_b\rightarrow0}\lim_{\ep_c\rightarrow0}}\frac{(-1)^{\ep_{1+j}}\ep_{1+j}+(-1)^{\ep_{2+j}}\ep_{2+j}}{\ep_{1+j}+\ep_{2+j}}\\
&\quad +\sum_{j=0}^2{ (-1)^{m_{3+j}}c(m_{1+j}, m_{2+j}; m_{3+j})}.
\end{align}
Since ${\displaystyle\lim_{\ep_a\rightarrow0}\lim_{\ep_b\rightarrow0}\lim_{\ep_c\rightarrow0}}\frac{(-1)^{\ep_{1+j}}\ep_{1+j}+(-1)^{\ep_{2+j}}\ep_{2+j}}{\ep_{1+j}+\ep_{2+j}}=1$, we only need to prove 
\begin{align}\label{eq:goal}
&\sum_{j=0}^2 (-1)^{m_{1+j}+m_{2+j}}b(m_{1+j}, m_{2+j}; m_{3+j})+\sum_{j=0}^2 (-1)^{m_{3+j}}c(m_{1+j}, m_{2+j}; m_{3+j})\\
&= \begin{cases} 1 & (m_1,m_2,m_3)=(0,0,0),\\ 0 & (m_1,m_2,m_3)\neq (0,0,0).\end{cases}
\end{align}
Since the left side of \eqref{eq:goal} coincides with the coefficient of $\frac{t_1^{m_1}t_2^{m_2}t_3^{m_3}}{m_1!m_2!m_3!}$ in the left side of \eqref{eq:gen_func}, the desired result follows from Lemma \ref{lem:gen}
\end{proof}

\section{Conclusion}

 By Theorem \ref{thm:main}, one can define the symmetric double omega values $\omega_{\mathcal{S}}(k_1,k_2,k_3)$ for non-positive integers $k_1,k_2,k_3\in \Z_{\le0}$ by
\begin{equation}\label{eq:limit_omega}
\omega_{\mathcal{S}}(k_1,k_2,k_3)=
{\lim_{\ep_a\rightarrow0}\lim_{\ep_b\rightarrow0}\lim_{\ep_c\rightarrow0}}\omega_{\mathcal{U}}(k_1+\ep_1,k_2+\ep_2,k_3+\ep_3)\mod \pi^2\mathcal{Z},
\end{equation}
which are either 0 or 1.
On the other hand, for $k_1,k_2,k_3\in \Z_{\le0}$ one computes
\begin{align}
\omega_{\mathcal{A}}(k_1,k_2,k_3)&=\left(\sum_{\substack{n_1+n_2<p\\n_1,n_2\ge1}} n_1^{-k_1}n_2^{-k_2}(p-n_1-n_2)^{-k_3}\mod p\right)_p\\
&=\left( \sum_{m_2=2}^{p-1}\sum_{m_1=1}^{m_2-1}m_1^{-k_1}(m_2-m_1)^{-k_2}(-m_2)^{-k_3}\mod p\right)_p\\
&=(-1)^{-k_2-k_3}\sum_{l=0}^{-k_2} (-1)^l \binom{-k_2}{l} \zeta_{\mathcal{A}}(k_1+k_2+l,k_3-l)
\end{align}
in $\mathcal{A}$.
Hence by \cite[Theorem 1.3]{Komori18}, for non-positive integers $k_1,k_2,k_3\in \Z_{\le0}$ we have
\[ \omega_{\mathcal{A}}(k_1,k_2,k_3)=\begin{cases} (1)_p & (k_1,k_2,k_3)=(0,0,0)\\ (0)_p & (k_1,k_2,k_3)\neq (0,0,0)\end{cases}.\]
Therefore the conjecture \eqref{eq:BTTConjecture} holds when we restrict $\boldsymbol{k}$ to elements in $\Z^3_{\le0}$.
Moreover, for integers $k_1,k_2,k_3$ such that $k_1\le k_2=0\le k_3$, as a consequence of \cite[Proposition 1.6]{Komori21}, we obtain the correspondence between the limit value \eqref{eq:limit_omega} and $\omega_{\mathcal{A}}(k_1,k_2,k_3)$, because we have $\omega_{\mathcal{U}}(k_1,0,k_3)=(-1)^{k_3}\zeta_{\mathcal{U}}(k_1,k_3)$ and $\omega_{\mathcal{A}}(k_1,0,k_3)=(-1)^{k_3}\zeta_{\mathcal{A}}(k_1,k_3)$.

Now a possible analogue of \eqref{eq:KomoriConjecture} for multiple zeta values of Mordell--Tornheim type is as follows.
Suppose that a certain limit ${\displaystyle\lim_{\boldsymbol{\ep}\rightarrow \boldsymbol{0}}}'\ \omega_{\mathcal{U}}(\boldsymbol{k}+\boldsymbol{\ep})$ exists in $\mathcal{Z}$ for all $\boldsymbol{k}\in \mathbb{I}$.
Then one can define the symmetric multiple omega values $\omega_{\mathcal{S}}(\boldsymbol{k})\in \mathcal{Z}/\pi^2\mathcal{Z}$ for all $\boldsymbol{k}\in \mathbb{I}$ by
\[\omega_{\mathcal{S}}(\boldsymbol{k})={\lim_{\boldsymbol{\ep}\rightarrow \boldsymbol{0}}}'\ \omega_{\mathcal{U}}(\boldsymbol{k}+\boldsymbol{\ep}) \mod \pi^2 \mathcal{Z}.\] 
One would ask if
\[\sum_{\boldsymbol{k}\in \mathbb{I}} a_{\boldsymbol{k}} \omega_{\mathcal{S}}(\boldsymbol{k})=0 \ \stackrel{?}{\Longleftrightarrow}  \ \sum_{\boldsymbol{k}\in \mathbb{I}} a_{\boldsymbol{k}} \omega_{\mathcal{A}}(\boldsymbol{k})=0\]
holds for a finite subset $\{a_{\boldsymbol{k}}\in\Q\mid \boldsymbol{k}\in \mathbb{I}, r\neq1\}$ of $\Q$.


\section*{Acknowledgments}
This work is partially supported by
JSPS KAKENHI Grant Number JP16H06336 and 20K14294.
{The authors thank Yasushi Komori for pointing out factual mistakes in the early paper}.


\end{document}